# Deflickering Vision-Based Occupancy Networks through Lightweight Spatio-Temporal Correlation

### v3

@@ Abstract
Index Terms:
vision-based occupancy network, deflickering, spatio-temporal correlation, plug-and-play.
@@ I Introduction
Vision-based Occupancy Networks (VONs) [1, 2, 3, 4, 5] have emerged as a powerful paradigm for reconstructing surrounding environments from ego-centric multi-view images. Typically, VONs represent the scene as a structured grid of voxels [6] by learning accurate mappings from 2D visual cues to 3D occupancy representations. This approach enables the extraction of both geometric structure and semantic understanding, thereby facilitating various applications, including indoor navigation [7, 8] and autonomous driving [9, 10, 11, 5].
@@
Despite these progresses, the occupancy results generated by existing VONs often exhibit inaccuracies and visual imperfections. A particularly problematic phenomenon, which we term flickering, manifests as instability in the constructed scenes: objects may appear or disappear abruptly, accompanied by various artifacts. This lack of temporal consistency and accuracy not only confuses the driving control system but also seriously degrades the visual experience for human drivers. The primary causes of flickering include sensor noise, occlusion, model limitations, and thresholding issues. As illustrated in Fig. 1, a pedestrian occluded by a tree provides incomplete visual cues, posing a significant challenge to the prediction models. In this case, SurroundOcc [12], which operates in a frame-by-frame “detection” paradigm, fails to detect the pedestrian in Frame #24. The temporary absence of the person will create a flickering effect when the results are displayed to the user.
Fig. 1: Enhancing existing VONs with OccLinker. (a) A standard 3D VON [12], which ignores historical data and thus misses small objects in certain frames; (b) A history-aware VON [13] that leverages historical frames, but still generates incomplete voxels due to suboptimal data association; (c) Our module, when integrated with the 3D VON [12], achieves smoother and more complete occupancy via more effective spatio-temporal correlation.
An effective approach to mitigate the temporal inconsistency problem is to incorporate historical information while predicting the current occupancy. Several recent methods [14, 13, 15] have explored this history-aware approach. While successful in reducing missing objects, these methods often incur significant computational costs. To balance accuracy and fast reconstruction, OPUS [14] employs a learnable query-based encoder-decoder architecture and retains coarse-grained concatenation of raw image features during early fusion stages. BEVDet4D-Occ [13] reduces redundancy by performing spatio-temporal alignment from a Bird’s Eye View (BEV) perspective. However, these methods aggregate spatio-temporal features and project them into the current latent space in a somewhat indiscriminate manner, leading to redundant information that may suppress object detection. As illustrated in the third column of Fig. 1, the pedestrian appears unclear in the BEVDet4D-Occ results, likely due to the coarse-grained integration of historical information during data fusion.
@@
Fig. 2: Comparison with history-aware VON methods in terms of mIoU performance and running costs. Points closer to the bottom-left indicate better efficiency, with lower memory usage and faster inference. Larger circles denote higher prediction quality with mIoU shown in numbers. Our method, when integrated with ViewFormer, achieves the best mIoU and the lowest memory consumption, while maintaining competitive inference latency.
@@
We present a novel method, called OccLinker, to address the aforementioned challenges. OccLinker is a plugin design that integrates with existing occupancy networks for improved performance. It achieves so by discriminating and extracting relevant motion-static attention for occupancy prediction, with a highlight on lightweight and effective spatio-temporal correlation. OccLinker employs a three-stage framework. First, it converts historical static and motion cues into queries through sparse tokenization. Next, an independent latent correlation module is introduced to attentively associate current features with historical cues within a shared latent space. Finally, the system generates correction occupancy that enables refinement of the base network’s results. Our method offers two key advantages:
@@
  • Seamless integration with state-of-the-art (SOTA) VONs for significantly enhanced prediction accuracy. When integrated with SurroundOcc [12] on its benchmark, it elevates IoU and mIoU by 5.18% and 1.82%, respectively. Similarly, when combined with ViewFormer [16] on the Occ3D [6] benchmark, it achieves accuracy gains of 0.34% in IoU and 2.08% in mIoU;
@@
  • Low cost through lightweight correlation operation, which surpasses recent history-aware VONs [14, 13, 17] with much lower computational costs. For instance, OPUS-L [14] requires 10.33 GB of GPU memory and 0.16 seconds per frame for inference. In contrast, ViewFormer+OccLinker only needs 4.58 GB memory and 0.12 seconds. Notably, the OccLinker plug-in itself uses just 105 MB of memory and adds only 0.01 seconds of latency.
@@
  • We devise a new temporal consistency metric to help analyze the effectiveness of various occupancy methods. Extensive experiments are conducted on both SurroundOcc [12] and Occ3D [6] benchmarks, showing the of advantages of our approach in accuracy while maintaining real-time inference efficiency.
@@ II Related Work
3D scene reconstruction technology has evolved rapidly in autonomous driving. Early studies focused on mathematical methods [18] and camera parameters [19, 20, 21] to facilitate 3D scene reconstruction from 2D inputs. MonoScene [18] explored the path of scene reconstruction from 2D to 3D using monocular RGB images, setting a groundbreaking benchmark. Later, more learning methods were applied. 1) Multi-view reconstruction [22, 23], which integrates multiple perspectives for enhanced depth estimation. 2) Multimodal methods combining RGB with depth [24, 25] or RGB with LiDAR [26] data, improving the robustness and precision of reconstructions.
@@
3D VONs extend the capabilities of 3D scene reconstruction technologies to enhance dynamic and precise environmental perception for autonomous vehicles. This task involves learning the associations between pixel-level visual cues and the occupancy states of 3D voxels. Recent approaches often employ Transformers [27] to effectively model these complex relationships, achieving significant advancements in accuracy and efficiency. 3D occupancy prediction have utilized multiple data inputs such as LiDAR [28, 29] and depth data [30, 31, 32, 33, 34, 35, 36] to predict voxel occupancy. However, with advancements in camera technology and image processing, current research predominantly adopts an image-based end-to-end approach [12, 37, 3, 38, 39]. These modern methods use data from single frames to predict occupancy, focusing on integrating rich semantic information at lower costs.
@@ II-C History-aware VONs
Recent advances in spatio-temporal perception have extended 3D VONs into temporal domains, primarily through two paradigms: (i) 4D VONs that estimate current and future states (e.g., Cam4DOcc [40], OccSora [41]), and (ii) history-aware VONs [17, 15, 13, 14] leverage historical features to refine current-frame predictions. In this paper, we focus primarily on the second category, where existing approaches may lead to spatio-temporal redundancy when fusing historical features [14] or fail to establish fine-grained correlations between pixel-level visual cues and occupied voxels [13]. To address this issue, we propose a solution by introducing a lightweight, plug-and-play module that effectively discriminates and extracts relevant motion-static attention, thereby complementing the current occupancy outcome.
@@ III-A Problem Formulation
In the context of vision-based occupancy networks (VONs) [9, 3, 12, 5, 11] for autonomous driving, the ego-vehicle typically collects two types of data: (i) multi-view RGB frames from surround-view cameras (typically six), captured at a relatively low and fixed frequency – we refer to these as keyframes; (ii) higher-frequency motion information between keyframes, such as optical flow or frame differences. As illustrated in Fig. 3, given a temporal window $\mathcal{W}$ of length $L$, the input includes a sequence of $L$ historical keyframes denoted by $S_{t-L:t}$, one current keyframe $S_{t}$, and a corresponding sequence of $L$ motion cues $M_{t-L:t}$, where each motion cue spans a short interval $\Delta l$. Let $|V|$ denote the number of cameras. For simplicity, we merge all multi-view inputs into a unified notation and do not distinguish between individual camera views. Both RGB and motion frames are assumed to have the same spatial resolution.
@@
Typically, a base 3D occupancy network $\Phi$ first encodes the current keyframe $\Phi$ using a shared 2D encoder $\Phi_{{\rm SE}}$ to extract static appearance features $s_{t}$. These features are then lifted into 3D space via a 2D-to-3D decoder $\Phi_{{\rm SD}}$, resulting in an initial occupancy prediction $O_{t}^{\prime}$. That is:
$O_{t}^{\prime}=\Phi_{{\rm SD}}^{*}(s_{t}:=\Phi_{{\rm SE}}^{*}(S_{t})),$ (1)
where $\cdot^{*}$ indicates frozen weights from the pre-trained foundational network.
Fig. 3: Overview of OccLinker. A frozen base 3D VON [12, 18, 16] runs at a fixed frequency to extract static texture features by $\Phi_{\rm SE}^{*}$ and generate initial occupancy prediction by $\Phi_{\rm SD}^{*}$. We employ $\Phi_{\rm ME}$ for extracting motion features of intermediate frames, and OccLinker uses the static and motion features from the recent frame window as queries, and the current static feature as the key and value. Through cross-attention [27], it constructs spatio-temporal correlations and outputs a correction term $\Delta O_{t}$ to refine the initial prediction.
@@
Our goal is to design a lightweight plug-in module, parameterized by $\Theta$, that leverages the motion and temporal context within the window $S_{t-L:t}$ and $M_{t-L:t}$ to enhance the foundational prediction $O_{t}^{\prime}$. To achieve this, the intermediate motion information is first encoded into a latent representation using a motion encoder $\Phi_{\rm ME}$, yielding temporal motion features denoted as $m_{t-L:t}$. Notably, both $s_{t}$ and $m_{t}$ share the same dimensionality of $|V|\times c\times w\times h$. Subsequently, a lightweight tokenization encoder $\Theta_{\rm TE}$ transforms both the static features $s_{t-L:t}$ and the motion features $m_{t-L:t}$ into a sparse token set. Finally, the module $\Theta$ establishes spatio-temporal correlations in the latent space to improve prediction smoothness and accuracy, while also compensating for missing voxels in the current occupancy estimate. The enhancement component $\Delta O_{t}$ is defined as:
$\Delta O_{t}=\Theta(s_{t},s_{t-L:t-1},m_{t-L:t-1}).$ (2)
@@
Notably, our method focuses on enhancing the current-frame occupancy using historical and motion cues, rather than forecasting future occupancy. The design principle emphasizes lightweight integration without retraining or modifying the base network.
III-B Overview
Fig. 3 presents the overall design of our OccLinker. In contrast to previous approaches [11, 16, 17, 15, 13, 14] that depend on redundant historical data, OccLinker compute the enhancement component $\Delta O_{t}$ using a dual cross-attention mechanism [27], designed to distill essential cues from both static and motion features while capturing compact latent correlations. First, we apply a sparse tokenization strategy to transform $s_{t-L:t}$ and $m_{t-L:t}$ into two separate sets of queries, as described in Section III-C. Then, we establish spatio-temporal correlation in a compact shared latent space, detailed in Section III-D. Finally, the optimization procedure is described in Section III-E, and the pseudo-code for the OccLinker inference pipeline is summarized in Algorithm 1.
III-C Sparse tokenization for query construction
Static feature encoder
We extract static spatial features from both the current and previous keyframes by directly reusing the 2D vision backbone modules, denoted as $\Phi_{\rm SE}$, from foundational vision-based occupancy networks (e.g., SurroundOcc [12] or ViewFormer [16]). Let $S_{t}$ represent the current surround-view image and $S_{t-k}$ denote the $k$-th previous keyframe within a temporal window of size $L$. The corresponding static features are computed as:
$s_{t}$ $=\Phi_{\rm SE}^{*}(S_{t}),$ (4)
$s_{t-L:t-1}$ $={\rm Concat}(\{\Phi_{\rm SE}^{*}(S_{t-k}):k\in[1,\cdots,L]\}),$
where $\Phi_{\rm SE}^{*}$ shares parameters across all time steps, and $\mathrm{Concat}(\cdot)$ denotes concatenation along the temporal dimension.
Motion feature encoder
Although occupancy prediction is typically performed using keyframes sampled at a fixed rate, the ego-vehicle continuously captures high-frequency RGB images. These intermediate frames, located between consecutive keyframes, contain rich motion cues that are valuable for spatial-temporal perception. To efficiently extract motion information, we adopt a simple yet effective frame-difference technique [42]. For each pair of adjacent keyframes $(S_{i},S_{i+1})$ within the historical window $[t-L,t]$, we compute the difference between any two intermediate frames indexed by $(\tau_{a},\tau_{b})$ satisfying $\tau_{a}<\tau_{b}$:
$F_{(i,i+1)}=S^{\tau_{b}}_{(i,i+1)}-S^{\tau_{a}}_{(i,i+1)}.$ (5)
$M_{t-L:t}=\mathrm{Concat}({F_{(i,i+1)}:i\in[t-L,t-1]}).$ (6)
Input: $S_{t-L:t}$, $M_{t-L:t}$, $\Phi_{\rm SE}$, $\Phi_{\rm SD}$, $\Phi_{\rm ME}$, $\Theta_{\rm TE}$, and $\Theta$.
Output: $\hat{O}_{t}$.
@@ III-C Sparse tokenization for query construction
3       Extract static features: $s_{t}\leftarrow\Phi_{\rm SE}^{*}(S_{t})$;
4       Predict initial occupancy result: $O_{t}^{\prime}\leftarrow\Phi_{\rm SD}^{*}(s_{t})$;
5       Extract motion features: $m_{t}\leftarrow\Phi_{\rm ME}(M_{t})$;
6       Append $(s_{t},m_{t})$ to $\mathcal{W}$ and discard the oldest entry if $|\mathcal{W}|>L$;
7       Retrieve $s_{t-L:t-1}$, $m_{t-L:t-1}$, and $s_{t}$ from $\mathcal{W}$;
8       Convert the above features into sparse tokens via Eq. (8);

### v4

@@ I Introduction
Vision-based Occupancy Networks (VONs) [OccFormer, zhang2025occnerf, VoxFormer, mei2024camera, yang2024adaptiveocc, jiang2024symphonize] have emerged as a powerful paradigm for reconstructing surrounding environments from ego-centric multi-view images. Typically, VONs represent the scene as a structured grid of voxels [Occ3D] by learning accurate mappings from 2D visual cues to 3D occupancy representations. This approach enables the extraction of both geometric structure and semantic understanding, thereby facilitating various applications, including indoor navigation [liu2024volumetric, li2023stereovoxelnet] and autonomous driving [xu2025exploiting, xu2025temporal, ouyang2024linkocc, yang2024adaptiveocc].
@@
Despite these progresses, the occupancy results generated by existing VONs often exhibit inaccuracies and visual imperfections. A particularly problematic phenomenon, which we term flickering, manifests as instability in the constructed scenes: objects may appear or disappear abruptly, accompanied by various artifacts. This lack of temporal consistency and accuracy not only confuses the driving control system but also seriously degrades the visual experience for human drivers. The primary causes of flickering include sensor noise, occlusion, model limitations, and thresholding issues. As illustrated in Fig. 1, a pedestrian occluded by a tree provides incomplete visual cues, posing a significant challenge to the prediction models. In this case, SurroundOcc [surroundOcc], which operates in a frame-by-frame “detection” paradigm, fails to detect the pedestrian in Frame #24. The temporary absence of the person will create a flickering effect when the results are displayed to the user.
Fig. 1: Enhancing existing VONs with OccLinker. (i) In the upper part, the first row shows three consecutive images where the ego-vehicle moves forward along the right side of the road. The blue box highlights a pedestrian as a sparse visual cue, and the second row provides the corresponding zoomed-in views. (ii) In the lower part, we compare different types of VON methods. The first row shows a standard 3D VON [surroundOcc] that ignores historical information, resulting in the pedestrian being missed in certain frames. The second row presents a history-aware VON [bevdet4d] that utilizes past frames but still produces incomplete voxels due to suboptimal temporal association. In contrast, our module, when integrated with the 3D VON [surroundOcc], achieves smoother and more complete occupancy predictions through effective spatio-temporal correlation.
An effective approach to mitigate the temporal inconsistency problem is to incorporate historical information when predicting the current occupancy. Recent history-aware VONs [opus, bevdet4d, SparseOcc_Liu, viewformer] have explored this direction and can be broadly categorized into two classes: dense Bird’s Eye View (BEV)/voxel-based temporal fusion methods and sparse query or representation based methods. Specifically:
  • Dense methods [bevdet4d, fb_occ, panoocc, viewformer, chen2025rethinking, DriveWorld, ye2024cvt] perform dense temporal modeling, typically by propagating, aligning, or aggregating historical BEV features or volumetric representations into the current frame. These approaches incur substantial computational and memory costs, often limiting the feasible temporal window or spatial resolution. For example, as shown in Fig. 2, BEVDet4D-Occ produces the largest inference latency. Moreover, dense temporal aggregation may propagate redundant or noisy features across broad spatial regions, leading to unstable predictions in dynamic areas. As illustrated in the lower part of Fig. 1, the pedestrian appears unclear in the BEVDet4D-Occ results, likely due to coarse-grained temporal fusion.
  • Sparse methods seek to reduce the complexity of dense voxel fusion by operating on sparse queries [opus, oh20253d] or representations [SparseOcc_Liu, liao2025stcocc, jiang2025gausstr]. By focusing computation on most relevant latent regions, methods such as SparseOcc and the OPUS series typically achieve lower inference latency (see Fig. 2). These methods often require multi-stage recovery or decoding processes to reconstruct dense occupancy. For instance, SparseOcc employs diffusion-based propagation combined with feature pyramids to densify non-empty regions across scales, whereas OPUS adopts a coarse-to-fine cross-attention mechanism to progressively decode sparse queries into 3D space. These designs increase memory usage and produce of reduced voxel-level accuracy.
@@
Fig. 2: Comparison with history-aware VON methods in terms of mIoU and running costs. Points closer to the bottom-left indicate better efficiency, with lower memory usage and faster inference. Larger circles denote higher prediction quality with mIoU shown in numbers. Our method, when integrated with ViewFormer, achieves the best mIoU and the lowest memory consumption, while maintaining competitive inference latency.
Different from existing methods, we explore an alternative approach: enhancing temporal consistency through a lightweight plug-in module that leverages historical information without redesigning the backbone, altering the fusion strategy, or retraining the base VON from scratch. Specifically, our approach only requires training a lightweight module on top of existing VON predictions or intermediate features, making it both practical and broadly applicable.
@@
To this end, we propose a novel method, termed OccLinker. Our approach improves occupancy prediction by discriminating and extracting relevant motion and static cues from historical frames, emphasizing lightweight yet effective spatio-temporal correlation modeling. OccLinker follows a three-stage framework. First, it converts historical static and motion cues into queries through sparse tokenization. Next, an independent latent correlation module is introduced to attentively associate current features with historical cues within a shared latent space. Finally, the system generates correction occupancy that enables refinement of the base network’s results. Our method offers two key advantages:
@@
  • Seamless integration with state-of-the-art (SOTA) VONs for significantly enhanced prediction accuracy. When integrated with SurroundOcc [surroundOcc] on its benchmark, it elevates IoU and temporal consistency of moving objects (as defined in the Experiment section) by 1.63% and 2.31, respectively. Similarly, when combined with ViewFormer [viewformer] on the Occ3D [Occ3D] benchmark, it achieves gains of 0.24% in IoU and 2.87 in temporal consistency;
@@
  • Low cost through lightweight correlation operation, which surpasses recent history-aware VONs [opus, bevdet4d, fb_occ] with much lower computational costs. For instance, OPUS-L [opus] requires 10.33 GB of GPU memory and 0.16 seconds per frame for inference. In contrast, ViewFormer+OccLinker only needs 4.58 GB memory and 0.12 seconds. Notably, the OccLinker plug-in itself uses just 105 MB of memory and adds only 0.01 seconds of latency.
@@
  • We devise a new temporal consistency metric to help analyze the effectiveness of various occupancy methods. Extensive experiments are conducted on both SurroundOcc [surroundOcc] and Occ3D [Occ3D] benchmarks, showing the of advantages of our approach in accuracy while maintaining real-time inference efficiency.
@@ II Related Work
3D scene reconstruction technology has evolved rapidly in autonomous driving. Early studies focused on mathematical methods [MonoScene] and camera parameters [3D_reconstruction_camera_para_1, 3D_reconstruction_camera_para_2, 3D_reconstruction_camera_para_3] to facilitate 3D scene reconstruction from 2D inputs. MonoScene [MonoScene] explored the path of scene reconstruction from 2D to 3D using monocular RGB images, setting a groundbreaking benchmark. Later, more learning methods were applied. 1) Multi-view reconstruction [Multi_View_reconstruction_1, Multi_View_reconstruction_2, shi2024occupancy], which integrates multiple perspectives for enhanced depth estimation. 2) Multimodal methods combining RGB with depth [multimodal_rgb_depth_1, multimodal_rgb_depth_2] or RGB with LiDAR [multimodal_rgb_lidar_1] data, improving the robustness and precision of reconstructions.
Fig. 3: Overview of OccLinker. A frozen base 3D VON [surroundOcc, MonoScene, viewformer] runs at a fixed frequency to extract static texture features by $\Phi_{\rm SE}^{*}$ and generate initial occupancy prediction by $\Phi_{\rm SD}^{*}$. We employ $\Phi_{\rm ME}$ for extracting motion features of intermediate frame differences, and OccLinker uses the static and motion features from the recent temporal window as queries, and the current static feature as the key and value. Through lightweight tokenization encoder $\phi$ and dual cross-attention module $\Theta$, it constructs spatio-temporal correlations and outputs a correction term $\Delta O_{t}$ to refine the initial prediction.
@@
3D VONs extend the capabilities of 3D scene reconstruction technologies to enhance dynamic and precise environmental perception for autonomous vehicles. This task involves learning the associations between pixel-level visual cues and the occupancy states of 3D voxels. Recent approaches often employ Transformers [attention_is_all_you_need] to effectively model these complex relationships, achieving significant advancements in accuracy and efficiency. 3D occupancy prediction have utilized multiple data inputs such as LiDAR [Panoptic_PolarNet, PointOcc, li2025occmamba] and depth data [S3CNet_depth, Li_2019_CVPR_depth, Li_2020_CVPR_depth, See_and_think_depth, Li_Zou_Li_Zhao_Gao_2020_depth, Depth_Based_Semantic_Scene_depth, Two_Stream_3D_Semantic_Scene_Completion_depth] to predict voxel occupancy. However, with advancements in camera technology and image processing, current research predominantly adopts an image-based end-to-end approach [surroundOcc, TPVFormer, VoxFormer, selfocc, RenderOcc]. These modern methods use data from single frames to predict occupancy, focusing on integrating rich semantic information at lower costs. However, due to the high-dimensional visual input and dense grid-based output, end-to-end methods suffer from significant computational intensity. To address this, more recent works [jiang2025gausstr, huang2024gaussianformer, huang2025gaussianformer, gan2025gaussianocc, zhu2025voxelsplat] have incorporated Gaussian point clouds as an intermediate representation. These point clouds are projected into the 3D occupancy space using splatting techniques, which significantly reduce memory consumption.
@@ II-C History-aware VONs
Recent advances in spatio-temporal perception have extended 3D VONs into temporal domains, primarily through two paradigms: (i) 4D VONs that estimate current and future states (e.g., Cam4DOcc [Cam4docc], OccSora [OccSora]), and (ii) history-aware VONs [fb_occ, SparseOcc_Liu, bevdet4d, opus] leverage historical features to refine current-frame predictions. In this paper, we focus primarily on the second category, which, as discussed in the Introduction, can be further classified into dense BEV or voxel-based temporal fusion [bevdet4d, fb_occ, panoocc, viewformer, chen2025rethinking, DriveWorld, ye2024cvt] and sparse query [opus, oh20253d] or representation-based methods [SparseOcc_Liu, liao2025stcocc, jiang2025gausstr]. For instance, the sparse method OPUS [opus] uses a learnable query-based encoder-decoder architecture and retains coarse-grained concatenation of raw image features during early fusion stages, while the dense method BEVDet4D-Occ [bevdet4d] performs spatio-temporal alignment from a BEV perspective. In contrast, we propose a different approach, using a lightweight plugin that effectively discriminates and extracts relevant motion-static attention, complementing the current occupancy predictions. Our method does not conflict with existing history-aware approaches and can seamlessly integrate with them or other 3D VONs to enhance both occupancy accuracy and temporal consistency.
@@ III-A Problem Formulation
In the context of vision-based occupancy networks (VONs) [xu2025exploiting, VoxFormer, surroundOcc, yang2024adaptiveocc, ouyang2024linkocc] for autonomous driving, the ego-vehicle typically collects two types of data: (i) multi-view RGB frames from surround-view cameras (typically six), captured at a relatively low and fixed frequency – we refer to these as keyframes; (ii) higher-frequency motion information between keyframes, such as optical flow or frame differences.
As illustrated in Fig. 3, given a temporal window $\mathcal{W}$ of length $L$, the input includes a sequence of $L$ historical keyframes denoted by $S_{t-L:t-1}$, one current keyframe $S_{t}$, and a corresponding sequence of $L$ motion cues $M_{t-L:t-1}$, where each motion cue represents frame differences over a short interval $\Delta l$. Let $|V|$ denote the number of cameras. For simplicity, we merge all multi-view inputs into a unified notation and do not distinguish between individual camera views. Both RGB and motion frames are assumed to have the same spatial resolution.
@@
Typically, a base 3D VON $\Phi$ first encodes the current keyframe $S_{t}$ using a 2D encoder $\Phi_{{\rm SE}}$ to extract static appearance features $s_{t}$. These features are then lifted into 3D space via a 2D-to-3D decoder $\Phi_{{\rm SD}}$, resulting in an initial occupancy prediction $O_{t}^{\prime}$. That is:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{t}=\Phi_{{\rm SE}}^{*}(S_{t}),\quad O_{t}^{\prime}=\Phi_{{\rm SD}}^{*}(s_{t}),$} (1)
where $\cdot^{*}$ indicates frozen weights from the pre-trained foundational network. Since the base VON operates in a frame-by-frame manner, the historical static features $s_{t-L:t}$ extracted by the frozen static encoder $\Phi_{\rm SE}^{*}$ are pre-stored and reused for temporal correlation modeling.
Our goal is to design a lightweight plug-in module, parameterized by $\Theta$, that associates the motion features within a temporal window of length $L$ with the static features $s_{t-L:t}$ extracted by the base VON, thereby enhancing the foundational prediction $O_{t}^{\prime}$ at each timestep $t$. To achieve this, the intermediate frame differences, denoted by $M_{t-L:t-1}$, is first encoded into a latent representation using a motion encoder $\Phi_{\rm ME}$, yielding temporal motion features:
$m_{t-L:t-1}=\Phi_{\rm ME}(M_{t-L:t-1}).$ (2)
@@
To enable efficient feature linking, we employ a lightweight tokenization encoder $\phi$ that transforms both the static features $s_{t-L:t}$ and the motion features $m_{t-L:t-1}$ into a compact sparse token set. Subsequently, the cross-attention module $\Theta$ establishes spatio-temporal correlations within this latent token space, enhancing prediction smoothness and accuracy while compensating for missing voxels in the current occupancy estimate. The resulting enhancement component $\Delta O_{t}$ is formulated as:
$\Delta O_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Theta}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi}(s_{t-L:t}),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi}(m_{t-L:t-1})).$ (3)
@@
Notably, our method focuses on enhancing the current-frame occupancy using historical and motion cues, rather than forecasting future occupancy. The design principle emphasizes lightweight integration without retraining or modifying the base network. Fig. 3 presents the overall design of our OccLinker. In contrast to previous approaches [ouyang2024linkocc, viewformer, fb_occ, SparseOcc_Liu, bevdet4d, opus] that depend on redundant historical data, OccLinker compute the enhancement component $\Delta O_{t}$ using a dual cross-attention mechanism [attention_is_all_you_need], designed to distill essential cues from both static and motion features while capturing compact latent correlations. First, we apply a sparse tokenization strategy to transform $s_{t-L:t}$ and $m_{t-L:t}$ into two separate sets of queries, as described in Section III-B. Then, we establish spatio-temporal correlation in a compact shared latent space, detailed in Section III-C. Finally, the optimization procedure is described in Section III-D, and the pseudo-code for the OccLinker inference pipeline is summarized in Algorithm 1.
Input: Keyframes $S_{t-L:t}$ and motion cues $M_{t-L:t}$; Frozen static encoder $\Phi_{\rm SE}^{*}$ and decoder $\Phi_{\rm SD}^{*}$ from the base VON; Trainable motion encoder $\Phi_{\rm ME}$ and trainable components in OccLinker, including tokenization encoder $\phi$ and dual cross-attention module $\Theta$.
Output: Enhanced occupancy prediction $\hat{O}_{t}$.
@@
 // Stage 1: Base VON Inference
3   Extract static features from the current frame: $s_{t}\leftarrow\Phi_{\rm SE}^{*}(S_{t})$;
4   Predict initial occupancy logits: $O_{t}^{\prime}\leftarrow\Phi_{\rm SD}^{*}(s_{t})$;
 // Stage 2: Spatio-Temporal Linking
5   Extract motion features from the high-frequency motion cues: $m_{t}\leftarrow\Phi_{\rm ME}(M_{t})$;
6   Append $(s_{t},m_{t})$ to the temporal window $\mathcal{W}$ and discard the oldest entry if $|\mathcal{W}|>L$;
7   Retrieve recent static and motion features $s_{t-L:t-1}$, $m_{t-L:t-1}$, and $s_{t}$ from $\mathcal{W}$;
8   Convert all retrieved features into sparse tokens via Eq. (9);